OpenStudy (maddiie):

Write each equation in slope- intercept form. Then identify the slope and the y-intercept for each line. 2x-y=12 I answered it like this y=2x+12. Is it right?

OpenStudy (anonymous):

y=2x-12 slope=2 y-intercept=-12

OpenStudy (maddiie):

But why minus

OpenStudy (maddiie):

I thought you add that

OpenStudy (anonymous):

2x-y=12 add y on both sides: 2x=12+y subtract 12 from both sides: 2x-12=y this is same as: y=2x-12 clear?

OpenStudy (anonymous):

sorry! u did tat wrong!:)

OpenStudy (maddiie):

Ohh... IT'S OK. I suck at Algrebra. I just don't want the answers i need to learn it too you know...

OpenStudy (anonymous):

believe urself maddiie..u can surely win and you will..i dont want my fans to be disheartened..:)

OpenStudy (maddiie):

Thanks with you i will learn this.